Solve the following equation..& find out the general solutions  sin3x+cos2x=0

Grade:11

1 Answers

Harsh Patodia IIT Roorkee
6 years ago
sin3x+cos2x=0
3sinx-4sin3x+1-2sin2x=0 (basic formulas)
(sinx+1)(4sin2x-2sinx-1)=0
sinx=-1 that is x=n pie+(-1)n3pie/2
and sinx=1\pm(5)½/4 that gives sin18 &sin72 so other two solutions will be x=npie+(-1)npie/10 and x=npie+(-1)n2pie/5